An asset with a book value of $80,000 is now being sold for $45,000. If the tax rate is 40%, what are the net proceeds from selling this asset?
Answer #1

Dear student thankyou for using homeworklib.

Please find the answer below.

Statement showing computation:

Capital gains tax credit= (sales value-book value)*tax rate

Capital gain tax credit= (45000-80000)*0.40= -14000

Net proceeds= 45000+14000= 59000

Net proceeds= $59000